two dice arr simultaneously rolled find probillity to getting sum of at least 11



1
Expert's answer
2021-06-29T03:41:24-0400

There are "6 \times 6 = 36" ways two dice could fall



There 3 ways (5+6, 6+5, 6+6) to get sum of at least 11

The probability to getting sum of at least 11

"P = \frac{3}{36}= \frac{1}{12}"